OpenStudy (anonymous):

SOMEHOW, I need to simplify this equation. -_- I don't think it's even possible... |Simplifying Algebraic Equations| 44^2 = 50^2 – 2 * 30 * 4 + 4^2

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

:D Hi tkhunny :D

OpenStudy (tkhunny):

44^2 = 50^2 – 2 * 30 * 4 + 4^2 Why not tackle it one piece at a time? 44^2 = 50^2 – (2 * 30 * 4) + 4^2 44^2 = ?? 50^2 = ?? 2*30*4 = ?? 4^2 = 16 Build it and simplify it. Of course, it's unlikely to work out very well.

OpenStudy (anonymous):

:OOO my hero.

OpenStudy (anonymous):

Thanks, tkhunny. I was confused D:

OpenStudy (anonymous):

44^2 = 44^2 50^2 = 50^2 60*4 = 240 4^2 = 16 44^2 = 50^2 - 240 + 16

OpenStudy (anonymous):

Am I on the right track, @tkhunny ?

OpenStudy (anonymous):

44^2 = 1936 50^2 = 2500

OpenStudy (anonymous):

1936 = 2500 - 240 + 16

OpenStudy (anonymous):

@jim_thompson5910 can you finish where tkhunny left off? He's offline now.

OpenStudy (ashleyisakitty):

44^2 = 50^2 – (2 * 30 * 4) + 4^2 1693 = 2500-250+16. Your work is correct so far. Now, it's time to simplify the left side. Can you do that?

OpenStudy (anonymous):

I shall try :D

OpenStudy (ashleyisakitty):

I mean the right side.

OpenStudy (anonymous):

I know, :D

OpenStudy (ashleyisakitty):

lol :P let me know what you get.

OpenStudy (anonymous):

2276

OpenStudy (anonymous):

1936 ≠ 2276 ?

OpenStudy (anonymous):

Awh, but I did the not equal to sign. The equal sign with the slash through it.

OpenStudy (ashleyisakitty):

Right. 1693 = 2276. That is, obviously, false. Therefore this problem is simply rubbish! This equation is, as you stated in your title, not possible. Nice work.

OpenStudy (anonymous):

yaaay :D